Question : Z, N×N, and Q are countable infinite. Where Z is integer, N is natural number and Q is ration number
Answer #1

In to show, ZA IN XIN and Q are countable infinite, e, we have to show, is there exist an injection from in two ways Z, INXINEach A; is countable infinite. P being the union of countable infinite number of countable infinite sets iss countable... PisCase-lo- min mtn .: 8(m, n) = 4 (mth-2) +n £4 (min-2) + H+ (m-1) = 4 (min-1) <4 (mith-2 min <mth L 7mtha mth-1 <4(mith